The length of a rectange is 6 inches less than twice its width. Find the dimension of the rectangle if its ares is 108 square inches.
Thank you

Answer by fractalier(1829)

Let the width be w. The length is then 2w - 6. Multiply them to get the area.
A = LW
A = (2w - 6)w
108 = w(2w - 6)
2w^2 - 6w - 108 = 0
w^2 - 3w - 54 = 0
(w - 9)(w + 6) = 0
w = 9 or w = -6 (impossible) so
w = 9 and
length = 2(9) - 6 = 12